# Measuring Vision-Language STEM Skills of Neural Models

### v1

@@ Appendix D Summary of Skills
Subject: Math Skill: select-figures-with-a-given-area Description: Which shape has an area of 7 square units? The shapes are made of unit squares. Picture: None Choices: [Uncaptioned image][Uncaptioned image] Answer index: 1
@@
Subject: Math Skill: select-fractions-equivalent-to-whole-numbers-using-models Description: Count the equal parts. What fraction does this picture show? Picture: [Uncaptioned image] Choices: [2/4, 4/8, 8/2, 2/8, ] Answer index: 2
Subject: Math Skill: select-solid-shapes Description: Which shape is a cone? Picture: None Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image] Answer index: 2
Subject: Math Skill: select-three-dimensional-shapes Description: Which shape is a rectangular prism? Picture: None Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image] Answer index: 2
Subject: Math Skill: select-two-dimensional-shapes Description: Which shape is a hexagon? Picture: None Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image] Answer index: 2
Subject: Math Skill: shapes-of-everyday-objects Description: Which is shaped like a cylinder? Picture: None Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image] Answer index: 1
@@
Subject: Science Skill: animal-adaptations-feet-and-limbs Description: Star-nosed moles are found in many parts of North America. They live in burrows. The moles eat earthworms and nuts, which they find in the soil. The feet of the star-nosed mole are adapted for digging.Which animal’s feet are also adapted for digging? Picture: [Uncaptioned image] Choices: [Uncaptioned image][Uncaptioned image] Answer index: 0
@@
Subject: Science Skill: animal-adaptations-skins-and-body-coverings Description: Emerald tree boas live in the forests of South America. The tree boa is adapted to be camouflaged among green leaves.Which animal is also adapted to be camouflaged among green leaves? Picture: [Uncaptioned image] Choices: [Uncaptioned image][Uncaptioned image] Answer index: 1
Subject: Science Skill: classify-elementary-substances-and-compounds-using-models Description: Complete the statement.Nitrogen is The model below represents a molecule of nitrogen. Nitrogen gas makes up nearly 80% of the air you breathe. Picture: [Uncaptioned image] Choices: [an elementary substance, a compound, ] Answer index: 0
@@
Subject: Science Skill: classify-fruits-and-vegetables-as-plant-parts Description: People use lettuce plants for food. We usually eat the part of this plant that makes most of the food for the plant.Hint: A plant’s leaves make food. A plant’s seeds can grow into a new plant. Which part of the lettuce plant do we usually eat? Picture: [Uncaptioned image] Choices: [the leaves, the seeds, ] Answer index: 0
Subject: Science Skill: classify-matter-as-solid-liquid-or-gas Description: Is the water from a faucet a solid, a liquid, or a gas? Picture: [Uncaptioned image] Choices: [a solid, a liquid, a gas, ] Answer index: 1
Subject: Science Skill: classify-matter-as-solid-or-liquid Description: Is a coin a solid or a liquid? Picture: [Uncaptioned image] Choices: [a liquid, a solid, ] Answer index: 1
Subject: Science Skill: classify-rocks-and-minerals-by-color-and-shape Description: Select the black mineral. Picture: None Choices: [Uncaptioned image][Uncaptioned image] Answer index: 0
Subject: Science Skill: classify-rocks-and-minerals-by-color-shape-and-texture Description: Select the brown rock. Picture: None Choices: [Uncaptioned image][Uncaptioned image] Answer index: 1
@@
Subject: Science Skill: classify-rocks-as-igneous-sedimentary-or-metamorphic Description: Diorite is a type of rock. When melted rock cools below the earth’s surface, it can form diorite. Diorite is usually made of large mineral grains.What type of rock is diorite? Picture: [Uncaptioned image] Choices: [sedimentary, igneous, ] Answer index: 1
@@
Subject: Science Skill: classify-symbiotic-relationships Description: Which type of relationship is formed when an Alcon blue caterpillar lives in a Myrmica ant nest?Read the passage. Then answer the question. Alcon blue butterflies spend the first part of their lives as caterpillars that live with Myrmica ants. When a caterpillar lives with the ants, it mimics, or pretends to be, an ant. The caterpillar can mimic the ants by copying their smell. The caterpillar can also make noises that make it sound like a queen ant. Queen ants receive more food and better protection than any other ants in the nest. So, when the caterpillar mimics an ant, the ants feed and protect the caterpillar instead of other ants in the nest. Picture: [Uncaptioned image] Choices: [mutualistic, commensal, parasitic, ] Answer index: 2
@@
Subject: Science Skill: compare-ages-of-fossils-in-a-rock-sequence Description: This diagram shows fossils in an undisturbed sedimentary rock sequence.Which of the following fossils is older? Select the more likely answer. Picture: [Uncaptioned image] Choices: [Uncaptioned image][Uncaptioned image] Answer index: 1
Subject: Science Skill: compare-amplitudes-and-wavelengths-of-waves Description: Select the wave with the greater amplitude. Picture: None Choices: [Uncaptioned image][Uncaptioned image] Answer index: 0
@@
Subject: Science Skill: compare-amplitudes-wavelengths-and-frequencies-of-waves Description: Select the graph of the wave with the greater amplitude.The graphs below describe two waves. The waves are traveling at the same speed. Picture: None Choices: [Uncaptioned image][Uncaptioned image] Answer index: 1
@@
Subject: Science Skill: compare-ancient-and-modern-organisms-use-observations-to-support-a-hypothesis Description: Which statement supports the following hypothesis?The American lobster and Homarus hakelensis have similar adaptations to survive underwater. The American lobster and Homarus hakelensis have similar adaptations to survive underwater. Picture: [Uncaptioned image] Choices: [Homarus hakelensis used its claws to find food underwater, The American lobster uses its claws to find food underwater, The American lobster uses its claws to find food underwater, ] Answer index: 2
@@
Subject: Science Skill: compare-concentrations-of-solutions Description: The diagram below is a model of two solutions. Each green ball represents one particle of solute.Which solution has a higher concentration of green particles? Picture: [Uncaptioned image] Choices: [Solution B, Solution A, neither; their concentrations are the same, ] Answer index: 0
@@
Subject: Science Skill: compare-magnitudes-of-magnetic-forces Description: Think about the magnetic force between the magnets in each pair. Which of the following statements is true?The images below show two pairs of magnets. The magnets in different pairs do not affect each other. All the magnets shown are made of the same material, but some of them are different shapes. Picture: [Uncaptioned image] Choices: [The magnitude of the magnetic force is smaller in Pair 2., The magnitude of the magnetic force is the same in both pairs., The magnitude of the magnetic force is smaller in Pair 1., ] Answer index: 1
Subject: Science Skill: compare-properties-of-materials Description: Which is harder? Picture: None Choices: [Uncaptioned image][Uncaptioned image] Answer index: 1
@@
Subject: Science Skill: compare-properties-of-objects Description: Which property do these four objects have in common?Select the best answer. Picture: [Uncaptioned image] Choices: [sticky, sour, soft, ] Answer index: 2
@@
Subject: Science Skill: use-data-to-describe-world-climates Description: Which statement best describes the climate of Santa Fe?Hint: Summers in the Northern Hemisphere occur in June, July, and August. Winters in the Northern Hemisphere occur in December, January, and February. Santa Fe, New Mexico, is a city in the United States. It has a semiarid climate. Picture: [Uncaptioned image] Choices: [Winters have much lower temperatures than summers., Winters have less precipitation than summers., ] Answer index: 0
@@
Subject: Science Skill: use-evidence-to-classify-animals Description: Placental mammals are a group of animals with similar traits. The following traits can be used to identify placental mammals:They give birth to live offspring. They have fur or hair. Observe the animals and read the descriptions. Select the one animal that has all of the placental mammal traits listed above. Sea otters have very thick fur. Their fur helps keep them warm in cold water. Female sea otters give birth to live offspring in the water. Red salamanders do not have lungs! They can breathe through their moist, smooth skin. Adult red salamanders live near rivers or ponds. They lay eggs with no shells under rocks or logs. The baby red salamanders live underwater. Picture: None Choices: [Uncaptioned image][Uncaptioned image] Answer index: 0
@@
Subject: Science Skill: use-evidence-to-classify-mammals-birds-fish-reptiles-and-amphibians Description: Fish are a group of animals with similar traits. The following traits can be used to identify fish:They have fins, not limbs. They make eggs with no shells. Observe the animals and read the descriptions. Select the one animal that has all of the fish traits listed above. Thresher sharks hatch from eggs with no shells. They have a long tail and fins. They can use their tail to hit and stun their prey. Thresher sharks live in salt water. Greater flameback woodpeckers have feathers and two wings. They use their strong beaks to make holes in trees. The woodpeckers use these holes as nests for their eggs, which have white shells. Picture: None Choices: [Uncaptioned image][Uncaptioned image] Answer index: 0
@@
Subject: Science Skill: use-punnett-squares-to-calculate-probabilities-of-offspring-types Description: In a group of dachshund dogs, some individuals have rough fur and others have soft fur. In this group, the gene for the fur texture trait has two alleles. The allele for rough fur (F) is dominant over the allele for soft fur (f).This Punnett square shows a cross between two dachshund dogs. What is the probability that a dachshund dog produced by this cross will be homozygous dominant for the fur texture gene? Picture: [Uncaptioned image] Choices: [3/4, 0/4, 2/4, 1/4, 4/4, ] Answer index: 2
@@
Subject: Science Skill: use-punnett-squares-to-calculate-ratios-of-offspring-types Description: In a group of Syrian hamsters, some individuals have short fur and others have long fur. In this group, the gene for the fur length trait has two alleles. The allele for short fur (F) is dominant over the allele for long fur (f).This Punnett square shows a cross between two Syrian hamsters. What is the expected ratio of offspring with long fur to offspring with short fur? Choose the most likely ratio. Picture: [Uncaptioned image] Choices: [3:1, 1:3, 0:4, 2:2, 4:0, ] Answer index: 2
@@
Subject: Science Skill: use-scientific-names-to-classify-organisms Description: This organism is a mantled howler. Its scientific name is Alouatta palliata.Select the organism in the same species as the mantled howler. Picture: [Uncaptioned image] Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image] Answer index: 1
@@
Subject: Science Skill: weather-and-climate-around-the-world Description: Does this passage describe the weather or the climate?Hint: Weather is what the atmosphere is like at a certain place and time. Climate is the pattern of weather in a certain place. A cloud forest is a mountain ecosystem that is home to a wide variety of species. The skies were mostly clear last week over this cloud forest, which is in Ecuador. Picture: [Uncaptioned image] Choices: [weather, climate, ] Answer index: 0
@@
Subject: Technology Skill: cables Description: What kind of computer related plug or port do you see here? Picture: [Uncaptioned image] Choices: [USB type-A port, HDMI plug, VGA port, USB type-C plug, ] Answer index: 3
Subject: Technology Skill: font Description: Identify this font type Picture: [Uncaptioned image] Choices: [Times Ancient Roman, Matisse ITC, Human521 BT, Bookman Old Style, ] Answer index: 2
Subject: Technology Skill: icons Description: This is a(n old) logo of which famous app or program? Picture: [Uncaptioned image] Choices: [Acrobat Reader, Google Pay, Microsoft Office PowerPoint, GoFundMe, ] Answer index: 2
@@
Subject: Technology Skill: logo Description: This is (part of) a (former) logo of which computer related brand? Picture: [Uncaptioned image] Choices: [Imation, Cisco, Nintendo, Verbatim, ] Answer index: 3
Subject: Technology Skill: others Description: What is the function of this key? Picture: [Uncaptioned image] Choices: [Copy, Undo, Delete, Paste, ] Answer index: 1
Subject: Technology Skill: parts Description: What kind of computer component do you see here? Picture: [Uncaptioned image] Choices: [Power Supply Unit, Computer Fan, CPU Socket, Molex Connector, ] Answer index: 2
Subject: Technology Skill: peripherals Description: What kind of computer peripheral do you see here? Picture: [Uncaptioned image] Choices: [Floppy Disk, DVD Spindle, Tablet, Bluetooth Headset, ] Answer index: 3
Subject: Technology Skill: photo Description: What type of video game console do you see here? Picture: [Uncaptioned image] Choices: [Nintendo Wii, Microsoft Xbox One, Microsoft Xbox, Mattel Intellivision, ] Answer index: 0
Subject: Technology Skill: web Description: What meaning or function is usually associated with this web interface symbol? Picture: [Uncaptioned image] Choices: [Storage for deleted files, Computer games, Reload/Refresh, Send e-mail, ] Answer index: 2

### v2

@@ Appendix D Summary of Skills
Description: Which shape has an area of 7 square units? The shapes are made of unit squares.
Picture: None
Choices: [Uncaptioned image][Uncaptioned image]
Answer index: 1
Subject: Math
Skill: select-fractions-equivalent-to-whole-numbers-using-models
@@
Description: Count the equal parts. What fraction does this picture show?
Picture: [Uncaptioned image]
Choices: [2/4, 4/8, 8/2, 2/8, ]
Answer index: 2
Subject: Math
Skill: select-solid-shapes
Description: Which shape is a cone?
Picture: None
Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image]
Answer index: 2
Subject: Math
Skill: select-three-dimensional-shapes
Description: Which shape is a rectangular prism?
Picture: None
Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image]
Answer index: 2
Subject: Math
Skill: select-two-dimensional-shapes
Description: Which shape is a hexagon?
Picture: None
Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image]
Answer index: 2
Subject: Math
Skill: shapes-of-everyday-objects
Description: Which is shaped like a cylinder?
Picture: None
Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
Answer index: 1
@@
Subject: Science
Skill: animal-adaptations-feet-and-limbs
@@
Description: Star-nosed moles are found in many parts of North America. They live in burrows. The moles eat earthworms and nuts, which they find in the soil. The feet of the star-nosed mole are adapted for digging.Which animal’s feet are also adapted for digging?
Picture: [Uncaptioned image]
Choices: [Uncaptioned image][Uncaptioned image]
Answer index: 0
Subject: Science
Skill: animal-adaptations-skins-and-body-coverings
@@
Description: Emerald tree boas live in the forests of South America. The tree boa is adapted to be camouflaged among green leaves.Which animal is also adapted to be camouflaged among green leaves?
Picture: [Uncaptioned image]
Choices: [Uncaptioned image][Uncaptioned image]
Answer index: 1
Subject: Science
Skill: classify-elementary-substances-and-compounds-using-models
Description: Complete the statement.Nitrogen is The model below represents a molecule of nitrogen. Nitrogen gas makes up nearly 80% of the air you breathe.
Picture: [Uncaptioned image]
Choices: [an elementary substance, a compound, ]
Answer index: 0
Subject: Science
Skill: classify-fruits-and-vegetables-as-plant-parts
@@
Description: People use lettuce plants for food. We usually eat the part of this plant that makes most of the food for the plant.Hint: A plant’s leaves make food. A plant’s seeds can grow into a new plant. Which part of the lettuce plant do we usually eat?
Picture: [Uncaptioned image]
Choices: [the leaves, the seeds, ]
Answer index: 0
Subject: Science
Skill: classify-matter-as-solid-liquid-or-gas
Description: Is the water from a faucet a solid, a liquid, or a gas?
Picture: [Uncaptioned image]
Choices: [a solid, a liquid, a gas, ]
Answer index: 1
Subject: Science
Skill: classify-matter-as-solid-or-liquid
Description: Is a coin a solid or a liquid?
Picture: [Uncaptioned image]
Choices: [a liquid, a solid, ]
Answer index: 1
Subject: Science
Skill: classify-rocks-and-minerals-by-color-and-shape
Description: Select the black mineral.
Picture: None
Choices: [Uncaptioned image][Uncaptioned image]
Answer index: 0
Subject: Science
Skill: classify-rocks-and-minerals-by-color-shape-and-texture
Description: Select the brown rock.
Picture: None
Choices: [Uncaptioned image][Uncaptioned image]
Answer index: 1
Subject: Science
Skill: classify-rocks-as-igneous-sedimentary-or-metamorphic
@@
Description: Diorite is a type of rock. When melted rock cools below the earth’s surface, it can form diorite. Diorite is usually made of large mineral grains.What type of rock is diorite?
Picture: [Uncaptioned image]
Choices: [sedimentary, igneous, ]
Answer index: 1
Subject: Science
Skill: classify-symbiotic-relationships
@@
Description: Which type of relationship is formed when an Alcon blue caterpillar lives in a Myrmica ant nest?Read the passage. Then answer the question. Alcon blue butterflies spend the first part of their lives as caterpillars that live with Myrmica ants. When a caterpillar lives with the ants, it mimics, or pretends to be, an ant. The caterpillar can mimic the ants by copying their smell. The caterpillar can also make noises that make it sound like a queen ant. Queen ants receive more food and better protection than any other ants in the nest. So, when the caterpillar mimics an ant, the ants feed and protect the caterpillar instead of other ants in the nest.
Picture: [Uncaptioned image]
Choices: [mutualistic, commensal, parasitic, ]
Answer index: 2
Subject: Science
Skill: compare-ages-of-fossils-in-a-rock-sequence
@@
Description: This diagram shows fossils in an undisturbed sedimentary rock sequence.Which of the following fossils is older? Select the more likely answer.
Picture: [Uncaptioned image]
Choices: [Uncaptioned image][Uncaptioned image]
Answer index: 1
Subject: Science
Skill: compare-amplitudes-and-wavelengths-of-waves
Description: Select the wave with the greater amplitude.
Picture: None
Choices: [Uncaptioned image][Uncaptioned image]
Answer index: 0
Subject: Science
Skill: compare-amplitudes-wavelengths-and-frequencies-of-waves
@@
Description: Select the graph of the wave with the greater amplitude.The graphs below describe two waves. The waves are traveling at the same speed.
Picture: None
Choices: [Uncaptioned image][Uncaptioned image]
Answer index: 1
Subject: Science
Skill: compare-ancient-and-modern-organisms-use-observations-to-support-a-hypothesis
@@
Description: Which statement supports the following hypothesis?The American lobster and Homarus hakelensis have similar adaptations to survive underwater. The American lobster and Homarus hakelensis have similar adaptations to survive underwater.
Picture: [Uncaptioned image]
Choices: [Homarus hakelensis used its claws to find food underwater, The American lobster uses its claws to find food underwater, The American lobster uses its claws to find food underwater, ]
Answer index: 2
Subject: Science
Skill: compare-concentrations-of-solutions
@@
Description: The diagram below is a model of two solutions. Each green ball represents one particle of solute.Which solution has a higher concentration of green particles?
Picture: [Uncaptioned image]
Choices: [Solution B, Solution A, neither; their concentrations are the same, ]
Answer index: 0
Subject: Science
Skill: compare-magnitudes-of-magnetic-forces
@@
Description: Think about the magnetic force between the magnets in each pair. Which of the following statements is true?The images below show two pairs of magnets. The magnets in different pairs do not affect each other. All the magnets shown are made of the same material, but some of them are different shapes.
Picture: [Uncaptioned image]
Choices: [The magnitude of the magnetic force is smaller in Pair 2., The magnitude of the magnetic force is the same in both pairs., The magnitude of the magnetic force is smaller in Pair 1., ]
Answer index: 1
Subject: Science
Skill: compare-properties-of-materials
Description: Which is harder?
Picture: None
Choices: [Uncaptioned image][Uncaptioned image]
Answer index: 1
Subject: Science
Skill: compare-properties-of-objects
@@
Description: Which property do these four objects have in common?Select the best answer.
Picture: [Uncaptioned image]
Choices: [sticky, sour, soft, ]
Answer index: 2
@@
Subject: Science
Skill: use-data-to-describe-world-climates
@@
Description: Which statement best describes the climate of Santa Fe?Hint: Summers in the Northern Hemisphere occur in June, July, and August. Winters in the Northern Hemisphere occur in December, January, and February. Santa Fe, New Mexico, is a city in the United States. It has a semiarid climate.
Picture: [Uncaptioned image]
Choices: [Winters have much lower temperatures than summers., Winters have less precipitation than summers., ]
Answer index: 0
Subject: Science
Skill: use-evidence-to-classify-animals
@@
Description: Placental mammals are a group of animals with similar traits. The following traits can be used to identify placental mammals:They give birth to live offspring. They have fur or hair. Observe the animals and read the descriptions. Select the one animal that has all of the placental mammal traits listed above. Sea otters have very thick fur. Their fur helps keep them warm in cold water. Female sea otters give birth to live offspring in the water. Red salamanders do not have lungs! They can breathe through their moist, smooth skin. Adult red salamanders live near rivers or ponds. They lay eggs with no shells under rocks or logs. The baby red salamanders live underwater.
Picture: None
Choices: [Uncaptioned image][Uncaptioned image]
Answer index: 0
Subject: Science
Skill: use-evidence-to-classify-mammals-birds-fish-reptiles-and-amphibians
@@
Description: Fish are a group of animals with similar traits. The following traits can be used to identify fish:They have fins, not limbs. They make eggs with no shells. Observe the animals and read the descriptions. Select the one animal that has all of the fish traits listed above. Thresher sharks hatch from eggs with no shells. They have a long tail and fins. They can use their tail to hit and stun their prey. Thresher sharks live in salt water. Greater flameback woodpeckers have feathers and two wings. They use their strong beaks to make holes in trees. The woodpeckers use these holes as nests for their eggs, which have white shells.
Picture: None
Choices: [Uncaptioned image][Uncaptioned image]
Answer index: 0
Subject: Science
Skill: use-punnett-squares-to-calculate-probabilities-of-offspring-types
@@
Description: In a group of dachshund dogs, some individuals have rough fur and others have soft fur. In this group, the gene for the fur texture trait has two alleles. The allele for rough fur (F) is dominant over the allele for soft fur (f).This Punnett square shows a cross between two dachshund dogs. What is the probability that a dachshund dog produced by this cross will be homozygous dominant for the fur texture gene?
Picture: [Uncaptioned image]
Choices: [3/4, 0/4, 2/4, 1/4, 4/4, ]
Answer index: 2
Subject: Science
Skill: use-punnett-squares-to-calculate-ratios-of-offspring-types
@@
Description: In a group of Syrian hamsters, some individuals have short fur and others have long fur. In this group, the gene for the fur length trait has two alleles. The allele for short fur (F) is dominant over the allele for long fur (f).This Punnett square shows a cross between two Syrian hamsters. What is the expected ratio of offspring with long fur to offspring with short fur? Choose the most likely ratio.
Picture: [Uncaptioned image]
Choices: [3:1, 1:3, 0:4, 2:2, 4:0, ]
Answer index: 2
Subject: Science
Skill: use-scientific-names-to-classify-organisms
@@
Description: This organism is a mantled howler. Its scientific name is Alouatta palliata.Select the organism in the same species as the mantled howler.
Picture: [Uncaptioned image]
Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image]
Answer index: 1
Subject: Science
Skill: weather-and-climate-around-the-world
@@
Description: Does this passage describe the weather or the climate?Hint: Weather is what the atmosphere is like at a certain place and time. Climate is the pattern of weather in a certain place. A cloud forest is a mountain ecosystem that is home to a wide variety of species. The skies were mostly clear last week over this cloud forest, which is in Ecuador.
Picture: [Uncaptioned image]
Choices: [weather, climate, ]
Answer index: 0
Subject: Technology
Skill: cables
@@
Description: What kind of computer related plug or port do you see here?
Picture: [Uncaptioned image]
Choices: [USB type-A port, HDMI plug, VGA port, USB type-C plug, ]
Answer index: 3
Subject: Technology
Skill: font
Description: Identify this font type
Picture: [Uncaptioned image]
Choices: [Times Ancient Roman, Matisse ITC, Human521 BT, Bookman Old Style, ]
Answer index: 2
Subject: Technology
Skill: icons
Description: This is a(n old) logo of which famous app or program?
Picture: [Uncaptioned image]
Choices: [Acrobat Reader, Google Pay, Microsoft Office PowerPoint, GoFundMe, ]
Answer index: 2
Subject: Technology
Skill: logo
@@
Description: This is (part of) a (former) logo of which computer related brand?
Picture: [Uncaptioned image]
Choices: [Imation, Cisco, Nintendo, Verbatim, ]
Answer index: 3
Subject: Technology
Skill: others
Description: What is the function of this key?
Picture: [Uncaptioned image]
Choices: [Copy, Undo, Delete, Paste, ]
Answer index: 1
Subject: Technology
Skill: parts
Description: What kind of computer component do you see here?
Picture: [Uncaptioned image]
Choices: [Power Supply Unit, Computer Fan, CPU Socket, Molex Connector, ]
Answer index: 2
Subject: Technology
Skill: peripherals
Description: What kind of computer peripheral do you see here?
Picture: [Uncaptioned image]
Choices: [Floppy Disk, DVD Spindle, Tablet, Bluetooth Headset, ]
Answer index: 3
Subject: Technology
Skill: photo
Description: What type of video game console do you see here?
Picture: [Uncaptioned image]
Choices: [Nintendo Wii, Microsoft Xbox One, Microsoft Xbox, Mattel Intellivision, ]
Answer index: 0
Subject: Technology
Skill: web
Description: What meaning or function is usually associated with this web interface symbol?
Picture: [Uncaptioned image]
Choices: [Storage for deleted files, Computer games, Reload/Refresh, Send e-mail, ]
Answer index: 2